OpenStudy (anonymous):

find y'(pi/4) if y(x)=(6sinx)^x y'(pi/4)=

OpenStudy (anonymous):

does y'= 6^x sin^x(x) (x cot(x)+log(6 sin(x)))

myininaya (myininaya):

since the exponent is a variable then I would use logarithm differentiation

OpenStudy (anonymous):

ok so the x goes to the front

OpenStudy (anonymous):

xln(6sinx)

myininaya (myininaya):

So you have something in this form: \[y=f^g \] => \[\ln(y)=g \ln(f) \text{ for } f>0\] Then differentiate both sides like so: \[\frac{y'}{y}=g' \cdot \ln(f)+g \cdot \frac{f'}{f} \\ \text{ applied product rule here along with some other rules of course }\]

myininaya (myininaya):

Then you solve that for y' by multiplying y on both sides giving you: \[y'=y(g' \ln(f)+g \frac{f'}{f})\]

OpenStudy (anonymous):

after that where does the pi/4 go?

myininaya (myininaya):

after you find y' you replace x with pi/4

myininaya (myininaya):

because y'(pi/4) tells us to do that

myininaya (myininaya):

y'(pi/4) means to find y' first and then to take y' and evaluate it for x=pi/4

OpenStudy (anonymous):

ok ill finish working through it

OpenStudy (anonymous):

ok so y'(pi/4)=ln(6sin(x))+x(6cos(x))/(6sin(x))

OpenStudy (anonymous):

idk i probably screwed it up

myininaya (myininaya):

that is't y'(pi/4) y'(pi/4) will be a constant looks like you found y'/y

OpenStudy (anonymous):

ok what next

myininaya (myininaya):

recall your objective is to find y'(pi/4) which means you need to find y'

myininaya (myininaya):

first

OpenStudy (anonymous):

6^x sin^x(x) (x cot(x)+log(6 sin(x)))

myininaya (myininaya):

\[\frac{y'}{y}=\ln(6 \sin(x))+\frac{x \cos(x)}{\sin(x)}\] to solve for y' you multiply y on both sides \[y'(x)=(6\sin(x))^x(x \cot(x)+\ln(6 \sin(x)))\]

myininaya (myininaya):

which is what you wrote if you mean that log to be natural log

myininaya (myininaya):

now plug in pi/4

myininaya (myininaya):

Don't be nervous about your answer if it doesn't turn out pretty

myininaya (myininaya):

unless you are able to approximate

myininaya (myininaya):

what i'm saying is the exact answer is not pretty at all

myininaya (myininaya):

and it can't be simplify to a more pretty form

myininaya (myininaya):

of course you can simplify some parts but in all what i mean is you won't get it to look much prettier than it already is which is pretty ugly

OpenStudy (anonymous):

yeah i understand. this answer is cringe worthy 2^(pi/8) 3^(pi/4) (pi/4+log(3 sqrt(2)))

OpenStudy (anonymous):

2^(pi/8) 3^(pi/4) (pi/4+log(3 sqrt(2)))

OpenStudy (anonymous):

can i get a head start with this problem. i want to work through this on my own. what's troubling me is x such that (0,pi/2)

OpenStudy (anonymous):

thanks a lot btw

myininaya (myininaya):

no problem

myininaya (myininaya):

i will look at your screenshot and try to untrouble you

myininaya (myininaya):

Oh. I think they restricted the domain of the function because recall tan is not defined everywhere

myininaya (myininaya):

This problem is same as last let y=x^(2tan(x)) take the same process we did before

myininaya (myininaya):

Well it isn't a thinking thing. It is exactly what they did, not thinking, but pure knowing. lol